OpenStudy (anonymous):

Take the integral from 0 to 2 of (x^2-4x+4)^(1/2)dx is: Answer: 2 Please show steps of how to get the answer.

OpenStudy (tkhunny):

Have you noticed that \(x^{2} - 4x + 4 = (x-2)^{2}\)?

OpenStudy (tkhunny):

Holy Cow!!! x-2 is positive on [0,2] \(\sqrt{x^{2} - 4x + 4} = \sqrt{(x-2)^{2}} = x-2\) That's quite a bit easier to integrate.

OpenStudy (amonoconnor):

...agreed. I didn't notice that simplification before:/

OpenStudy (anonymous):

I tried simplifying it but got -2 instead of 2. :/

OpenStudy (amonoconnor):

So, (1/2x^2 - 2x), and plugged in it's: -2. Good work emele613! :)

OpenStudy (anonymous):

But that's not the right answer. :(

OpenStudy (anonymous):

It's positive 2.

OpenStudy (tkhunny):

@amonoconnor That was a lot of work. Good work having the gumption to drag through it!!

OpenStudy (amonoconnor):

Thank you! That's shat I appreciate seeing, so I try and do it for others:)

OpenStudy (amonoconnor):

*what

OpenStudy (tkhunny):

Whoops!!! x-2 is strictly NEGATIVE on [0,2]. Must have been standing on my head! \(\sqrt{(x-2)^{2}} = -(x-2) = 2-x\;for\;x\in[0,2]\) Sorry about that.

OpenStudy (accessdenied):

Oh, I didn't even catch that! Nice one. :P

OpenStudy (tkhunny):

Once in a while I see it if someone badgers me enough. The ORIGINAL integrand is strictly positive, so it made no sense to get a negative result after simplification.

OpenStudy (anonymous):

Thank you everyone! :) I would've not been able to get that.

OpenStudy (tkhunny):

Well, @amonoconnor 's "Brute Force" method would not have needed this subtle consideration. Kudos again to @amonoconnor for dragging through all that algebra.